### v1

@@ 5 Full Implementation
A basic game $\Gamma=\left(\mathcal{A},\Theta,p,\boldsymbol{c}\right)$ is a potential game if for every $\theta\in\Theta$, there exists an open superset $\widetilde{\mathcal{Y}}$ of $\mathcal{Y}$ in $\mathbb{R}^{\mathcal{A}}$ and a continuously differentiable function $\Phi_{\theta}\colon\widetilde{\mathcal{Y}}\to\mathbb{R}$, called potential, such that, for every $a\in\mathcal{A}$ and $\boldsymbol{y}\in\mathcal{Y}$,
$\frac{\partial\Phi_{\theta}(\boldsymbol{y})}{\partial y_{a}}=c_{a}(\boldsymbol{y},\theta).$ (5.1)
@@
Every game with two actions is a potential game. An important class of potential games is the one of congestion games defined by: a finite set of resources $\mathcal{E}$; a set of actions $\mathcal{A}\subseteq 2^{\mathcal{E}}$; for each resource $e\in\mathcal{E}$, a continuous cost function $c_{e}\colon\mathbb{R}_{+}\times\Theta\to\mathbb{R}_{+}$. The cost of choosing action $a\in\mathcal{A}$ in state $\theta$ is the sum of the costs of all resources used in $a$, $c_{a}(\boldsymbol{y},\theta)\coloneqq\sum_{e\in a}c_{e}(x_{e},\theta)$, where $x_{e}\coloneqq\sum_{a\in\mathcal{A},a\ni e}y_{a}$ is the load on resource $e$ induced by the flow $\boldsymbol{y}$. The potential function is
$\Phi_{\theta}(\boldsymbol{y})=\sum_{e\in\mathcal{E}}\int_{0}^{x_{e}}c_{e}(t,\theta)\ \textup{d}t.$ (5.2)
@@
When for every state, the cost function of each resource is increasing, the potential is convex. If the set of actions is the set of single resources ($\mathcal{A}=\{\{e\},e\in\mathcal{E}\}$) and the cost functions are strictly increasing, the potential is strictly convex.
@@
Consider a basic game $\Gamma$. If the potential is convex for each state, then for every information structure $\mathcal{I}$, all BWE of $(\Gamma,\mathcal{I})$ have the same cost profile. If the potential is strictly convex for each state, then for every information structure $\mathcal{I}$ there exists a unique BWE outcome of $(\Gamma,\mathcal{I})$.
To prove Proposition 5.2, we first show that a BWE of $(\Gamma,\mathcal{I})$ is equivalent to a Wardrop equilibrium of an auxiliary game with complete information and multiple populations which is also a potential game. If the basic game has a convex potential in each state, then the potential of the auxiliary game is also convex. Uniqueness of Wardrop equilibrium costs of the auxiliary game follows from Koessler et al., (2024, Proposition 4). As a consequence, we get uniqueness of BWE cost profiles for $(\Gamma,\mathcal{I})$. If the basic game has a strictly convex potential in each state, then the potential of the auxiliary game is a strictly convex function of the interim total flow profiles $(\boldsymbol{y}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$. Uniqueness of the equilibrium interim total flow profile follows then from the fact that the Wardrop equilibrium of the auxiliary multi-population game with complete information are the minimizers of the potential (see Sandholm,, 2001, Proposition 3.1).
Proof of Proposition 5.2.
@@
$C_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}})\coloneqq\sum_{\theta,\boldsymbol{\tau}^{-k}}p(\theta)\pi\left(\tau^{k},\boldsymbol{\tau}^{-k}\mid\theta\right)c_{a}\left(\left(\sum_{j\in\mathcal{K}}y_{b}^{j}(\tau^{j})\right)_{b\in\mathcal{A}},\theta\right)=\operatorname{\mathsf{P}}(\tau^{k})c_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}}).$ (5.3)
@@
Suppose that in each state $\theta$ there is a potential $\Phi_{\theta}$. Then let
@@
where we use the notation $y_{a}(\boldsymbol{\tau})=\sum_{k\in\mathcal{K}}y_{a}^{k}(\tau^{k})$ and $\boldsymbol{y}(\boldsymbol{\tau})=\left(y_{a}(\boldsymbol{\tau})\right)_{a\in\mathcal{A}}.$ The function $\Phi_{\pi}$ is a potential for the auxiliary game: it is immediate to check that
$\frac{\partial\Phi_{\pi}(\widehat{\boldsymbol{y}})}{\partial y_{a}^{k}(\tau^{k})}=C_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}}).$ (5.5)
If $\Phi_{\theta}$ is strictly convex for each $\theta$, then $\Phi_{\pi}$ is a strictly convex function of the vector $(\boldsymbol{y}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$. By Sandholm, (2001, Proposition 3.1), the equilibrium flows are the minimizers of the potential function. Thus, there is a unique equilibrium vector $(\boldsymbol{y}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$, and therefore a unique equilibrium outcome, given by (2.7). If $\Phi_{\pi}$ is (weakly) convex, there is a unique equilibrium cost profile $(C_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}}))_{k,\tau^{k},a}$ from Koessler et al., (2024, Proposition 4), and thus a unique BWE cost profile from Eq. 5.3. ∎
The next theorem is a full implementation result for games with strictly convex potential. It shows that if $\mu$ is a BCWE with finite support and rational flows, then there exists a direct information structure such that $\mu$ is implemented as a unique Bayesian Wardrop equilibrium outcome. Further, the information structure is direct and the BWE is obedient. If the support of $\mu$ is not finite or if some flows in the support are not rational, then $\mu$ is close to the unique BWE outcome $\mu^{\ast}$ for some direct information structure $\mathcal{I}$.
@@
Let $\Gamma$ be a basic game that has a strictly convex potential in each state. Let $\mu$ be a BCWE and let $\psi\colon\mathcal{Y}\times\Theta\to\mathbb{R}$ be a continuous function. For all $\varepsilon>0$, there exists a direct information structure $\mathcal{I}$ such that the Bayesian Wardrop equilibrium outcome $\mu^{\ast}$ of $(\Gamma,\mathcal{I})$ is unique and
@@
In addition, if $\mu$ has finite support with rational flows, then we can choose $\varepsilon=0$ and the obedient outcome $\breve{\mu}_{\mathcal{I}}$ is the unique Bayesian Wardrop equilibrium outcome $\mu^{\ast}$.
As before, $\psi$ represents any possible cost function for the designer. Also, for every $L>0$, we can choose the same $\mathcal{I}$ for all $L$-Lipschitz $\psi$.
The structure of the proof is as follows. Consider the construction given in the proof of Theorem 4.1: We approximate $\mu$ by a distribution over a finite set of rational flows and construct a direct information structure such that the obedient outcome is a Bayesian Wardrop $\alpha$-equilibrium for some $\alpha>0$, and is close to $\mu$. The difficulty is that the obedient outcome need not be an exact Bayesian Wardrop equilibrium and that we don’t know how to construct this BWE explicitly. We show that any Bayesian Wardrop $\alpha$-equilibrium is an $\alpha$-minimizer of the potential of the game with the constructed information structure, and that all $\alpha$-minimizers are $\varepsilon$-close to each other for some $\varepsilon>0$. In particular, the unique BWE outcome $\mu^{\ast}$ is close to the obedient outcome and thus, also close to $\mu$. This reasoning is summarized in the following proposition.
@@
Let $\Gamma$ be a basic game that has a strictly convex potential in each state. Let $\mu$ be a BCWE and let $\psi\colon\mathcal{Y}\times\Theta\to\mathbb{R}$ be a continuous function. For all $\varepsilon>0$, there exists $\alpha>0$ and a direct information structure $\mathcal{I}$ such that the Bayesian Wardrop equilibrium outcome $\mu^{\ast}$ of $(\Gamma,\mathcal{I})$ is unique, and for all $\tilde{\mu}\in\operatorname{\mathsf{BWE_{\alpha}}}(\mathcal{I})$,
$\left\lvert\int\psi(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)-\int\psi(\boldsymbol{y},\theta)\ \textup{d}\tilde{\mu}(\boldsymbol{y}\mid\theta)\right\rvert\leq\varepsilon,\quad\forall\theta\in\Theta.$ (5.7)
@@
Proof of Proposition 5.4.
We take up the construction from the proof of Theorem 4.1: fix $\alpha>0$, assume that $\mu$ has finite support, and consider the information structure $\mathcal{I}$ constructed there with parameter $\eta>0$, such that the obedient flow profile $\breve{\boldsymbol{y}}_{\mathcal{I}}$ with outcome $\breve{\mu}_{\mathcal{I}}$ is a Bayesian Wardrop $\alpha$-equilibrium and
@@
Suppose that the potential is (weakly) convex in each state. For all $\alpha>0$, any Bayesian Wardrop $\alpha$-equilibrium of the game with information structure $\mathcal{I}$, is an $\alpha$-minimizer of the associated potential.
Proof of 5.5.
@@
$C_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}})\leq C_{b}^{k,\tau^{k}}(\widehat{\boldsymbol{y}})+\alpha\operatorname{\mathsf{P}}(\tau^{k})\leq C_{b}^{k,\tau^{k}}(\widehat{\boldsymbol{y}})+\alpha,$ (5.9)
where $C_{a}^{k,\tau^{k}}$ is defined as in (5.3). This implies that, for all $k\in\mathcal{K}$, for all $\tau^{k}\in\mathcal{T}^{k}$,
$\sum_{a\in\mathcal{A}}\frac{y_{a}^{k}(\tau^{k})}{\gamma^{k}}C_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}})\leq\min_{b\in\mathcal{A}}C_{b}^{k,\tau^{k}}(\widehat{\boldsymbol{y}})+\alpha=\min_{\widehat{\boldsymbol{z}}}\sum_{a\in\mathcal{A}}\frac{z_{a}^{k}(\tau^{k})}{\gamma^{k}}C_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}})+\alpha.$ (5.10)
@@
$\sum_{a\in\mathcal{A}}\left(y_{a}^{k}(\tau^{k})-z_{a}^{k}(\tau^{k})\right)C_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}})\leq\gamma^{k}\alpha.$ (5.11)
@@
$\left\langle\nabla\Phi_{\pi}(\widehat{\boldsymbol{y}}),\widehat{\boldsymbol{y}}-\widehat{\boldsymbol{z}}\right\rangle=\sum_{k}\sum_{\tau^{k}}\sum_{a\in\mathcal{A}}\left(y_{a}^{k}(\tau^{k})-z_{a}^{k}(\tau^{k})\right)C_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}})\leq\sum_{k}\gamma^{k}\alpha=\alpha,$ (5.12)
and since $\Phi_{\pi}$ is convex, this gives
$\Phi_{\pi}(\widehat{\boldsymbol{y}})-\Phi_{\pi}(\widehat{\boldsymbol{z}})\leq\left\langle\nabla\Phi_{\pi}(\widehat{\boldsymbol{y}}),\widehat{\boldsymbol{y}}-\widehat{\boldsymbol{z}}\right\rangle\leq\alpha.$ (5.13)
Thus $\Phi_{\pi}(\widehat{\boldsymbol{y}})\leq\min\Phi_{\pi}+\alpha$. ∎
From strict convexity, let $(\boldsymbol{y}^{*}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$ be the unique Bayesian Wardrop equilibrium flow vector. We argue now that for any $\alpha$-minimizer $\widehat{\boldsymbol{y}}$ of $\Phi_{\pi}$, the flow vector $(\boldsymbol{y}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$ is close to $(\boldsymbol{y}^{*}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$.

### v2

@@ 5.1 Potential Games
A basic game $\Gamma=\left(\mathcal{A},\Theta,p,\boldsymbol{u}\right)$ is called a potential game if for every $\theta\in\Theta$, there exists an open superset $\widetilde{\mathcal{Y}}$ of $\mathcal{Y}$ in $\mathbb{R}^{\mathcal{A}}$ and a continuously differentiable function $\Phi_{\theta}\colon\widetilde{\mathcal{Y}}\to\mathbb{R}$ such that, for every $a\in\mathcal{A}$ and $\boldsymbol{y}\in\mathcal{Y}$,
$\frac{\partial\Phi_{\theta}(\boldsymbol{y})}{\partial y_{a}}=u_{a}(\boldsymbol{y},\theta).$ (5.1)
The function $\Phi_{\theta}$ is called the potential in state $\theta$.
@@
As a first example, consider symmetric congestion games, which are defined by: a finite set of resources $\mathcal{E}$; a set of actions $\mathcal{A}\subseteq 2^{\mathcal{E}}$; for each resource $e\in\mathcal{E}$, a continuous cost function $c_{e}\colon\mathbb{R}_{+}\times\Theta\to\mathbb{R}_{+}$. The payoff of choosing action $a\in\mathcal{A}$ in state $\theta$ is the negative of the sum of the costs of all resources used in $a$,
$u_{a}(\boldsymbol{y},\theta)\coloneqq-\sum_{e\in a}c_{e}(x_{e},\theta),$ (5.2)
where $x_{e}\coloneqq\sum_{a\in\mathcal{A},a\ni e}y_{a}$ is the load on resource $e$ induced by the flow $\boldsymbol{y}$. The potential function of a congestion game is
$\Phi_{\theta}(\boldsymbol{y})=-\sum_{e\in\mathcal{E}}\int_{0}^{x_{e}}c_{e}(t,\theta)\ \textup{d}t.$ (5.3)
@@
When the cost function of each resource is increasing for every state, the potential is concave. In singleton congestion games each action uses a single resource, i.e., $\mathcal{A}=\{\{e\},e\in\mathcal{E}\}$. In this class of games, if the cost functions are strictly increasing, the potential is strictly concave.
The following lemma shows that, unlike what happens in games with a finite set of players (see Monderer and Shapley,, 1996), not all potential games are congestion games.
Lemma 5.2.
Fix an arbitrary state $\theta$. A game with a potential $\Phi_{\theta}\in C^{3}$ is a congestion game only if
$\frac{\partial^{3}\Phi_{\theta}(\boldsymbol{y})}{\partial y_{a}^{2}\partial y_{b}}=\frac{\partial^{3}\Phi_{\theta}(\boldsymbol{y})}{\partial y_{b}^{2}\partial y_{a}}.$ (5.4)
Proof.
Consider a potential game with a potential function $\Phi_{\theta}\in C^{3}$, and suppose that it is a congestion game, with cost functions $c_{e}(\,\cdot\,,\theta)$, which are then twice continuously differentiable. Combine Eqs.˜5.1 and 5.2 to get,
$\frac{\partial\Phi_{\theta}(\boldsymbol{y})}{\partial y_{a}}=-\sum_{e\in a}c_{e}(x_{e},\theta).$ (5.5)
This implies that
$\frac{\partial^{2}\Phi_{\theta}(\boldsymbol{y})}{\partial y_{a}\partial y_{b}}=-\sum_{e\in a\cap b}c^{\prime}_{e}(x_{e},\theta)=\frac{\partial^{2}\Phi_{\theta}(\boldsymbol{y})}{\partial y_{b}\partial y_{a}}.$ (5.6)
Observe that $x_{e}=\sum_{a^{\prime}\in\mathcal{A},a\ni e}y_{a^{\prime}}$. Because we consider only resources $e\in a\cap b$, the flow $y_{a}$ or $y_{b}$ only appears in the sum $y_{a}+y_{b}$. Therefore, taking derivative once more with respect to either $y_{a}$ or $y_{b}$ yields then the same result. ∎
To go beyond congestion games, we consider the class of aggregative games. This class was implicitly introduced by Dubey et al., (1980), formalized by Corchón, (1994), and recently studied by Lahkar, (2017) for nonatomic players and a finite set of actions. Aggregative games are also given as examples of potential games in Sandholm, (2010, chapter 3).
An aggregative game is defined as follows. Let the action set $\mathcal{A}$ be a finite subset of $\mathbb{R}_{+}$, so actions can be seen as quantities. For $y\in\Delta(\mathcal{A})$, let the aggregate quantity be $Q(\boldsymbol{y})=\sum_{a\in\mathcal{A}}ay_{a}$. Each payoff $u_{a}$ depends only on the chosen action $a$ and on the aggregate quantity, i.e., for each $a\in\mathcal{A}$ and for each $\theta\in\Theta$, there exists a function $g_{a}(\,\cdot\,,\theta)$ such that $u_{a}(y,\theta)=g_{a}(Q(\boldsymbol{y}),\theta)$.
Consider payoffs of the following form:
$\forall a\in\mathcal{A},u_{a}(y,\theta)=a\rho(Q(\boldsymbol{y}),\theta)-C_{a}(\theta),$ (5.7)
where $\rho(\,\cdot\,,\theta)$ is a function of the aggregate quantity $Q(\boldsymbol{y})$, and $C_{a}(\theta)$ is a number that does not depend on $\boldsymbol{y}$. This game admits the potential:
$\Phi_{\theta}(\boldsymbol{y})=\int_{0}^{Q(\boldsymbol{y})}\rho(t,\theta)\ \textup{d}t-\sum_{a}C_{a}(\theta)y_{a}.$ (5.8)
If $\rho(\,\cdot\,,\theta)$ is a decreasing function of the quantity, then the potential is concave. [^6] Sandholm, (2010) and Lahkar, (2017) also consider the case of increasing $\rho(\,\cdot\,,\theta)$, which models public good contribution games where the value of the good is increasing with the total contribution, and players incur contribution costs. In this case, the potential is convex. This potential can be interpreted as a demand function and the game is a Cournot game, where $C_{a}(\theta)$ is the production cost of quantity $a$. In this model, both the demand and the production costs may be uncertain. In Example˜3.2 only production costs are uncertain and the potential is strictly concave.
With the help of Lemma˜5.2, assuming that $\rho(\,\cdot\,,\theta)$ is twice continuously differentiable, we can check whether the game given by Eq.˜5.7 is a congestion game or not. Suppose that $\rho^{\prime\prime}(\,\cdot\,,\theta)>0$ and assume that there exist two distinct actions $a,b\in\mathcal{A}\setminus\{0\}$. Then
$\frac{\partial^{3}\Phi_{\theta}(\boldsymbol{y})}{\partial y_{a}^{2}\partial y_{b}}=a^{2}b\rho^{\prime\prime}(Q(\boldsymbol{y}),\theta)\neq ab^{2}\rho^{\prime\prime}(Q(\boldsymbol{y}),\theta)=\frac{\partial^{3}\Phi_{\theta}(\boldsymbol{y})}{\partial y_{b}^{2}\partial y_{a}}.$ (5.9)
Therefore, the game is not a congestion game. Take for instance a demand function given by $\rho(q,\theta)=\kappa_{\theta}/(\lambda_{\theta}+q)$.
@@
A potential game with a potential $\Phi_{\theta}\in C^{2}$ is aggregative if and only if Eq.˜5.7 holds.
Proof.
Take an aggregative game with payoffs $u_{a}(y,\theta)=g_{a}(Q(\boldsymbol{y}),\theta)$ and suppose it admits a potential $\Phi_{\theta}$. Then,
$\frac{\partial\Phi_{\theta}(\boldsymbol{y})}{\partial y_{a}}$ $=g_{a}(Q(\boldsymbol{y}),\theta)$ (5.10)
and, for all $a,b\in\mathcal{A}$,
$\frac{\partial^{2}\Phi_{\theta}(\boldsymbol{y})}{\partial y_{b}\partial y_{a}}$ $=bg^{\prime}_{a}(Q(\boldsymbol{y}),\theta)=ag^{\prime}_{b}(Q(\boldsymbol{y}),\theta)=\frac{\partial^{2}\Phi_{\theta}(\boldsymbol{y})}{\partial y_{a}\partial y_{b}}.$ (5.11)
So for any two actions $a,b\neq 0$, the two ratios
$\frac{g^{\prime}_{a}(Q(\boldsymbol{y}),\theta)}{a}=\frac{g^{\prime}_{b}(Q(\boldsymbol{y}),\theta)}{b}$ (5.12)
are equal to some common function of $Q(\boldsymbol{y})$. Call this function $\rho^{\prime}(Q(\boldsymbol{y}),\theta)$. Eq.˜5.7 then follows by integration, where $C_{a}(\theta)$ is an integration constant. Note that for action 0, Eq.˜5.11 implies that $g^{\prime}_{0}(Q(\boldsymbol{y}),\theta)=0$, so Eq.˜5.7 is valid for $a=0$ as well. ∎
Finally, notice that every game with two actions is a potential game, though the potential need not be concave (see, e.g., Example˜3.1).
5.2 Full Implementation
The next proposition shows that for every potential game with a (strictly) concave potential, the BWE payoff profile (outcome) of the game extended by an information structure is unique.
Proposition 5.4.
Consider a basic game $\Gamma$. If the potential is concave for each state, then for every information structure $\mathcal{I}$, all BWE of $(\Gamma,\mathcal{I})$ have the same payoff profile. If the potential is strictly concave for each state, then for every information structure $\mathcal{I}$ there exists a unique BWE outcome of $(\Gamma,\mathcal{I})$.
To prove Proposition˜5.4, we first show that a BWE of $(\Gamma,\mathcal{I})$ is equivalent to a Wardrop equilibrium of an auxiliary game with complete information and multiple populations, which is also a potential game. If the basic game has a concave potential in each state, then the potential of the auxiliary game is also concave. Uniqueness of Wardrop equilibrium payoffs in the auxiliary game follows from Koessler et al., (2025, Proposition 4). As a consequence, we get uniqueness of BWE payoff profiles for $(\Gamma,\mathcal{I})$. If the basic game has a strictly concave potential in each state, then the potential of the auxiliary game is a strictly concave function of the interim total flow profiles $(\boldsymbol{y}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$. Uniqueness of the equilibrium interim total flow profile follows from the fact that the Wardrop equilibrium of the auxiliary multi-population game with complete information are the maximizers of the potential (see Sandholm,, 2001, proposition 3.1).
Proof of Proposition˜5.4.
Consider a basic $\Gamma$ and an information structure $\mathcal{I}=\left(\boldsymbol{\gamma},\mathcal{T},\pi\right)$. Define a multi-population auxiliary game with complete information where populations are indexed by $(k,\tau^{k})$, for all $k\in\mathcal{K},\tau^{k}\in\mathcal{T}^{k}$. A flow profile for population $(k,\tau^{k})$ is $\boldsymbol{y}^{k}(\tau^{k})=\left(y_{a}^{k}(\tau^{k})\right)_{a\in\mathcal{A}}\in\Delta_{\gamma^{k}}(\mathcal{A})$. Given the interim flow profile $\widehat{\boldsymbol{y}}=\left(y^{k}_{a}(\tau^{k})\right)_{k,\tau^{k},a}$, the payoff function of population $(k,\tau^{k})$ is
@@ 5.2 Full Implementation
Suppose that in each state $\theta$ there is a potential $\Phi_{\theta}$. Then, let
@@
where we use the notation $y_{a}(\boldsymbol{\tau})=\sum_{k\in\mathcal{K}}y_{a}^{k}(\tau^{k})$ and $\boldsymbol{y}(\boldsymbol{\tau})=\left(y_{a}(\boldsymbol{\tau})\right)_{a\in\mathcal{A}}.$ The function $\Phi_{\pi}$ is a potential for the auxiliary game and
[…]
If $\Phi_{\theta}$ is strictly concave for each $\theta$, then $\Phi_{\pi}$ is a strictly concave function of the vector $(\boldsymbol{y}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$. By Sandholm, (2001, Proposition 3.1), the equilibrium flows are the maximizers of the potential function. Thus, there is a unique equilibrium vector $(\boldsymbol{y}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$, and therefore a unique equilibrium outcome, given by (2.7). If $\Phi_{\pi}$ is (weakly) concave, uniqueness of the equilibrium payoff profile $(U_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}}))_{k,\tau^{k},a}$ follows from Koessler et al., (2025, proposition 4). Uniqueness of the BWE payoff profile follows from Eq.˜5.13. ∎
The next theorem is a full implementation result for games with a strictly concave potential. It shows that if $\mu$ is a BCWE with finite support and rational flows, then there exists a direct information structure such that $\mu$ is implemented as a unique Bayesian Wardrop equilibrium outcome. Furthermore, the information structure is direct and the BWE is obedient. If the support of $\mu$ is not finite or if some flows in the support are not rational, then $\mu$ is close to the unique BWE outcome $\mu^{\ast}$ for some direct information structure $\mathcal{I}$.
@@
Let $\Gamma$ be a basic game with a strictly concave potential in each state. Let $\mu$ be a BCWE and let $\psi\colon\mathcal{Y}\times\Theta\to\mathbb{R}$ be a continuous function. For all $\varepsilon>0$, there exists a direct information structure $\mathcal{I}$ such that the Bayesian Wardrop equilibrium outcome $\mu^{\ast}$ of $(\Gamma,\mathcal{I})$ is unique and
@@
In addition, if $\mu$ has finite support with rational flows, then we can choose $\varepsilon=0$ and the unique Bayesian Wardrop equilibrium outcome $\mu^{\ast}$ is the obedient outcome $\breve{\mu}_{\mathcal{I}}$.
As before, $\psi$ represents any possible welfare function for the designer. Also, for every $L>0$, we can choose the same $\mathcal{I}$ for all $L$-Lipschitz $\psi$.
The structure of the proof is as follows. Consider the construction given in the proof of Theorem˜4.1: Approximate $\mu$ by a distribution over a finite set of rational flows and construct a direct information structure such that the obedient outcome is a Bayesian Wardrop $\alpha$-equilibrium for some $\alpha>0$, and is close to $\mu$. The difficulty is that the obedient outcome need not be an exact Bayesian Wardrop equilibrium and that we don’t know how to construct this BWE explicitly. We show that any Bayesian Wardrop $\alpha$-equilibrium is an $\alpha$-maximizer of the potential of the game with the constructed information structure, and that all $\alpha$-maximizers are $\varepsilon$-close to each other for some $\varepsilon>0$. In particular, the unique BWE outcome $\mu^{\ast}$ is close to the obedient outcome and thus, also close to $\mu$. This reasoning is summarized in the following proposition.
@@
Let $\Gamma$ be a basic game with a strictly concave potential in each state. Let $\mu$ be a BCWE and let $\psi\colon\mathcal{Y}\times\Theta\to\mathbb{R}$ be a continuous function. For all $\varepsilon>0$, there exist $\alpha>0$ and a direct information structure $\mathcal{I}$ such that the Bayesian Wardrop equilibrium outcome $\mu^{\ast}$ of $(\Gamma,\mathcal{I})$ is unique, and for all $\widetilde{\mu}\in\operatorname{\mathsf{BWE_{\alpha}}}(\mathcal{I})$,
$\left\lvert\int\psi(\boldsymbol{y},\theta)\ \textup{d}\mu(\boldsymbol{y}\mid\theta)-\int\psi(\boldsymbol{y},\theta)\ \textup{d}\widetilde{\mu}(\boldsymbol{y}\mid\theta)\right\rvert\leq\varepsilon,\quad\forall\theta\in\Theta.$ (5.17)
@@
Proof of Proposition˜5.6.
We take up the construction from the proof of Theorem˜4.1: fix $\alpha>0$, assume that $\mu$ has finite support, and consider the information structure $\mathcal{I}$ constructed there with parameter $\eta>0$, such that the obedient flow profile $\breve{\boldsymbol{y}}_{\mathcal{I}}$ with outcome $\breve{\mu}_{\mathcal{I}}$ is a Bayesian Wardrop $\alpha$-equilibrium and
@@
Suppose that the potential is (weakly) concave in each state. For all $\alpha>0$, any Bayesian Wardrop $\alpha$-equilibrium of the game with information structure $\mathcal{I}$, is an $\alpha$-maximizer of the associated potential.
Proof of ˜5.7.
@@
$U_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}})\geq U_{b}^{k,\tau^{k}}(\widehat{\boldsymbol{y}})-\alpha\operatorname{\mathsf{P}}(\tau^{k})\geq U_{b}^{k,\tau^{k}}(\widehat{\boldsymbol{y}})-\alpha,$ (5.19)
where $U_{a}^{k,\tau^{k}}$ is defined as in (5.13). This implies that, for all $k\in\mathcal{K}$, for all $\tau^{k}\in\mathcal{T}^{k}$,
$\sum_{a\in\mathcal{A}}\frac{y_{a}^{k}(\tau^{k})}{\gamma^{k}}U_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}})\geq\max_{b\in\mathcal{A}}U_{b}^{k,\tau^{k}}(\widehat{\boldsymbol{y}})-\alpha=\max_{\widehat{\boldsymbol{z}}}\sum_{a\in\mathcal{A}}\frac{z_{a}^{k}(\tau^{k})}{\gamma^{k}}U_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}})-\alpha.$ (5.20)
@@
$\sum_{a\in\mathcal{A}}\left(y_{a}^{k}(\tau^{k})-z_{a}^{k}(\tau^{k})\right)U_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}})\geq-\gamma^{k}\alpha.$ (5.21)
@@
$\left\langle\nabla\Phi_{\pi}(\widehat{\boldsymbol{y}}),\widehat{\boldsymbol{y}}-\widehat{\boldsymbol{z}}\right\rangle=\sum_{k}\sum_{\tau^{k}}\sum_{a\in\mathcal{A}}\left(y_{a}^{k}(\tau^{k})-z_{a}^{k}(\tau^{k})\right)U_{a}^{k,\tau^{k}}(\widehat{\boldsymbol{y}})\geq-\sum_{k}\gamma^{k}\alpha=-\alpha,$ (5.22)
and since $\Phi_{\pi}$ is concave, this gives
$\Phi_{\pi}(\widehat{\boldsymbol{y}})-\Phi_{\pi}(\widehat{\boldsymbol{z}})\geq\left\langle\nabla\Phi_{\pi}(\widehat{\boldsymbol{y}}),\widehat{\boldsymbol{y}}-\widehat{\boldsymbol{z}}\right\rangle\geq-\alpha.$ (5.23)
Thus $\Phi_{\pi}(\widehat{\boldsymbol{y}})\geq\max\Phi_{\pi}-\alpha$. That is, any Bayesian Wardrop $\alpha$-equilibrium is an $\alpha$-maximizer of the potential. ∎
From strict concavity, let $(\boldsymbol{y}^{*}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$ be the unique Bayesian Wardrop equilibrium flow vector. We argue now that for any $\alpha$-maximizer $\widehat{\boldsymbol{y}}$ of $\Phi_{\pi}$, the flow vector $(\boldsymbol{y}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$ is close to $(\boldsymbol{y}^{*}(\boldsymbol{\tau}))_{\boldsymbol{\tau}}$.